# Galton-Watson processes, simple varieties of trees and Khinchin families

### v5

@@ Abstract.
In this note, we introduce a unified analytic framework that connects Bienaymé–Galton–Watson processes, simple varieties of trees and Khinchin families. Using Lagrange’s inversion formula, we derive new coefficient-based expressions for extinction probabilities and reinterpret them as boundary phenomena tied to the domain of the inverse of the solution to Lagrange’s equation. This perspective reveals an additional link between combinatorial and probabilistic models, simplifying classical arguments and yielding new results. It also leads to a computationally efficient method for simulating Galton–Watson processes via power series coefficients.
@@ 1. Introduction
One of the main results of this note is the following new coefficient-based expression for the probability of extinction of a Galton–Watson process: let $\psi\in\mathcal{K}$ be a power series with radius of convergence $R_{\psi}>0$ and Khinchin family $(Y_{t})$ and let $Y_{t}$ be the offspring distribution of our process, then we have
@@
here $A_{n}$ is the $n$-th coefficient of the solution of Lagrange’s equation $g(z)=z\psi(g(z))$ with data $\psi\in\mathcal{K}$. These coefficients can be computed in terms of $\psi$ by using Lagrange’s inversion formula. We refer the reader to Theorem 2 below for a detailed statement and proof of this result.
A simple variety of trees is a family of weighted rooted plane trees in which each node receives a weight that depends on its number of children (outdegree). The weight of the whole tree is the product of the different node-weights, see [10, p. 452]. Throughout this note we choose those node-weights to be the coefficients of a power series $\psi\in\mathcal{K}$ (see eq. (5.1) below). For such a fixed $\psi$, the sequence $A_{n}$ records the sum of the weights over all rooted plane trees with $n$ nodes (see Proposition A below). Specific choices of $\psi$ yield concrete families of combinatorial trees; for instance, $\psi(z)=e^{z}$ correspond to rooted labelled Cayley trees, while $\psi(z)=1/(1-z)$ correspond to rooted plane trees.
@@
Here, we simultaneously consider all possible Galton–Watson processes with offspring distribution given by a Poisson random variable with parameter $t>0$. Notice that $q(t)\equiv 1$, for any $t\leq\tau=1$ (to wit, the series above is actually identical to 1) and also that the radius of convergence of $g(z)$ is given by $\rho=\tau/\psi(\tau)=1/e$, see Theorem 1 below. When we cross the radius of convergence of $g(z)$, that is, when $t>\tau=1$, the inversion relation ${t}/{\psi(t)}=g^{-1}(t)$ no longer holds, and consequently the extinction probability should satisfy the inequality $q(t)<1$. Observe that for $t>\tau=1$, the function $t/\psi(t)=te^{-t}$ is strictly decreasing, see Lemma 4.1 below.
After deriving formula (1.1), we analyze the probability of extinction function in detail, establishing its continuity, differentiability (except at one point) and characterizing its asymptotic behavior as the parameter $t\uparrow R_{\psi}$. An explicit upper bound on the extinction probability function is also derived. Moreover, we establish several precise relationships between Khinchin families and random variables arising in Galton–Watson processes. We identify some of these random variables as appropriately reparameterized Khinchin families. Beyond its theoretical interest, this framework offers an efficient method for simulating Galton-Watson processes by means of power series and the numerical computation of its coefficients. We believe that this perspective will be valuable to both combinatorialists and probabilists, as it provides an explicit bridge between these viewpoints.
@@ 1.1. Notations
[…]
@@ Acknowledgements
The author thanks Professor José L. Fernández for many valuable conversations.
@@ 2. Khinchin families
Notice that $f(t)>0$ for each $t\in[0,R)$. We define $X_{0}\equiv 0$. Formally, the definition of $X_{0}$ is consistent with the general expression for $t\in(0,R)$, with the convention that $10^{0}=$ 1, meaning that $\mathbf{P}(X_{0}=0)=1$.
@@ 2.2. The classes $\mathcal{K}_{s}$ and $\mathcal{K}^{\star}$
[…]
@@ 3.1. The combinatorial class of rooted plane trees
We denote by $(\mathcal{G},|\cdot|)$ the combinatorial class of rooted plane trees, where the size function ||$\cdot$|| ⋅ | gives the number of nodes in a given tree. This family of rooted trees is codified, in the language of the combinatorial classes, by means of the specification
@@ 3.3. Galton–Watson processes
with $Y_{j}$ i.i.d. copies of $Y$ and $Z_{0}=Y$. The process $(Z_{n})_{n\geq 0}$ is a Galton-Watson process. The mean of the offspring distribution $m=\mathbf{E}[Y]$ determines extinction: if $m\leq 1$, extinction occurs almost surely ($q=1$); if $m>1$, the extinction probability $q<1$. The probability $q$ is the minimal solution of the equation
@@ 4. Lagrange’s inversion formula
The following classical result describes an explicit relationship between the coefficients of the power series $\psi$ and the solution $g$, thereby providing an explicit power series representation of $g$ as a solution to Lagrange’s equation ()($\star$)( ⋆ ).
@@ 4.1. Function $t/\psi(t)$
$g(z/\psi(z))=z\,,\quad\mbox{ for }z\in\Omega_{g}=g(\mathbb{D}(0,\rho))\subseteq\mathbb{D}(0,\tau)\,.$
@@
that is, the image of the interval $[0,\rho]$ by $g$ is the interval $g([0,\tau/\psi(\tau)])=[0,\tau]$. One may verify this claim directly (using that $g$ is increasing and continuous on the closure of $\mathbb{D}(0,\rho)$), or else deduce it as a corollary of Theorem B below.
@@
    is strictly increasing on the interval $[0,\tau)$
@@ 4.2. Radius of convergence of $g$
In case that $\psi\in\mathcal{K}^{\star}$ we have $R_{g}=\rho=\tau/\psi(\tau)$
@@
Note that when $\psi\in\mathcal{K}^{\star}$, the Khinchin family $(Z_{t})$ associated with the solution of Lagrange’s equation is naturally parametrized by $t\in[0,\rho]$, where $\rho=\tau/\psi(\tau)$. For further details about extending Khinchin families families to boundary values we refer to [9].
@@ 5. Lagrange’s equation and simple varieties of trees
Fix $\psi\in\mathcal{K}$ with radius of convergence $R_{\psi}>0$. A simple variety of trees is a family of weighted rooted plane trees whose weighted generating function is the solution of Lagrange’s equation with data $\psi$, that is, is a power series $g$ verifying the relation
@@
Here $k_{j}(a)$ counts the number of nodes in $a$ with exactly $j$ descendants. The next result is standard, see [23] for similar arguments.
@@
Let $\psi\in\mathcal{K}$ a power series with radius of convergence $R_{\psi}>0$, then
@@ 6. The parametric Galton Watson process
Given a power series $\psi\in\mathcal{K}$ and Khinchin family $(Y_{t})$, in this section we study the Galton-Watson process with offspring distribution $Y_{t}$. This construction gives rise to a family of parametric Galton-Watson process $T_{t}$ which depends on $t\in(0,R_{\psi})$. For each $t\in(0,R_{\psi})$ the rooted random tree $T_{t}$ takes values on the Polish space $\mathbb{T}=\mathcal{G}\cup\mathcal{G}_{\infty}$.
@@ 6.1. The Galton-Watson process with offspring distribution $Y_{t}$
with the convention that $10^{0}=$ 1. Recall that $k_{j}(a)$ counts the number of nodes in $a$ with exactly $j$ descendants. This measure assigns probability 00 to the singletons in $\mathcal{G}_{\infty}$.
@@
with the convention that $10^{0}=$ 1. Here, $\omega(a)$ denotes the weight of the rooted plane tree $a\in\mathcal{G}_{n}$.
@@
these equations express that the tree has $n$ nodes and $n-1$ edges (or, equivalently, $n-1$ nodes that descend from a parent node, we don’t count the root), we obtain that
@@ 6.2. Probability generating functions
The power series $g_{t}(z)$ verifies Lagrange’s equation with data $\psi_{t}$ by definition. For any $t\in(0,R)$, denote $f_{t}(z)={g(tz/\psi(t))}/{t}$. Notice that
@@ 6.3. Extinction probability
The probability of extinction $q(t)$ is continuous on the interval $[0,R_{\psi})$.
@@
If $\psi\in\mathcal{K}\setminus\mathcal{K}^{\star}$, we have $q(t)\equiv 1$, for any $t\in[0,R_{\psi})$, which is continuous. Assume that $\psi\in\mathcal{K}^{\star}$, the function $t/\psi(t)$ is continuous for any $t\in[0,R_{\psi})$. Now, combining Theorem 1 with Lemma 4.1, we find that
@@
$q(t)\sim\frac{\psi(0)}{\psi(t)}\,,\quad\text{ as }t\uparrow R_{\psi}\,.$
@@
Theorem 2 gives that
@@
  • If $R_{\psi}<\infty$, then $\lim_{t\uparrow R_{\psi}}\psi(t)<\infty$ and the probability $q(t)$ can be extended to $t=R_{\psi}$.
@@
Now we give an upper estimate for the extinction probability.
@@
For any $t\in(0,R_{\psi})\setminus\{\tau\}$, combining Theorem 2, equation (7.1) and the identity $\psi(tq(t))=q(t)\psi(t)$ we obtain that
@@ 8.1. The uniform probability
with the convention that $10^{0}=$ 1. By means of $\omega$, see Proposition A above, we can build all the families of simple varieties of trees.
@@
again with the convention that $10^{0}=$ 1. Assume, for simplicity, that $A_{n}\neq 0$ for any $n\geq 1$, otherwise, we can restrict $n\geq 1$ to values satisfying $n\equiv 1\mod Q_{\psi}$. First, we consider the measure on $\mathcal{G}$ with weights $\omega_{\mathbf{P}}$ and $\omega$. For a rooted plane tree $a\in\mathcal{G}$ we compute its measure as follows

### v7

@@ Abstract.
In this note, we introduce a unified analytic framework that connects Bienaymé–Galton–Watson processes, simple varieties of trees and Khinchin families. Using Lagrange’s inversion formula, we derive new coefficient-based expressions for extinction probabilities and reinterpret them as boundary phenomena tied to the domain of the inverse of the solution to Lagrange’s equation. This perspective reveals an additional link between combinatorial and probabilistic models, simplifying classical arguments and yielding new results. It also leads to a computationally efficient method for simulating probabilities associated to Galton–Watson processes via power series coefficients.
@@ 1. Introduction
One of the main results of this note is the following new coefficient-based expression for the probability of extinction of a Galton–Watson process: let $\psi\in\mathcal{K}$ be a power series with radius of convergence $R_{\psi}>0$ and Khinchin family $(Y_{t})$ and let $Y_{t}$ be the offspring distribution of our process, then the probability of extinction $q(t)$ is given by:
@@
here $A_{n}$ is the $n$-th coefficient of the solution of Lagrange’s equation $g(z)=z\psi(g(z))$ with data $\psi\in\mathcal{K}$. These coefficients can be computed in terms of $\psi$ by using Lagrange’s inversion formula. We refer the reader to Theorem 6.4 below for a detailed statement and proof of this result.
A simple variety of trees is a family of weighted rooted plane trees in which each node receives a weight that depends on the number of its children (outdegree). The weight of the whole tree is the product of the different node-weights, see [10, p. 452]. Throughout this note we choose those node-weights to be the coefficients of a power series $\psi\in\mathcal{K}$ (see eq. (5.1) below). For such a fixed $\psi$, the sequence $A_{n}$ records the sum of the weights over all rooted plane trees with $n$ nodes (see Proposition A below). Specific choices of $\psi$ yield concrete families of combinatorial trees; for instance, $\psi(z)=e^{z}$ corresponds to rooted labelled Cayley trees, while $\psi(z)=1/(1-z)$ corresponds to rooted plane trees.
@@
Here, we simultaneously consider all possible Galton–Watson processes with offspring distribution given by a Poisson random variable with parameter $t>0$. Notice that $q(t)\equiv 1$, for any $t\leq\tau=1$ (to wit, the series above is actually equal to 1 for $t\leq 1$) and also that the radius of convergence of $g(z)$ is given by $\rho=\tau/\psi(\tau)=1/e$, see Theorem 4.2 below. When we cross the radius of convergence of $g(z)$, that is, when $t>\tau=1$, the inversion relation ${t}/{\psi(t)}=g^{-1}(t)$ no longer holds, and consequently the extinction probability should satisfy the inequality $q(t)<1$. Observe that for $t>\tau=1$, the function $t/\psi(t)=te^{-t}$ is strictly decreasing, see Lemma 4.1 below.
After deriving formula (1.1), we analyze the probability of extinction function in detail, establishing its continuity, differentiability (except at one point) and characterizing its asymptotic behavior as the parameter $t\uparrow R_{\psi}$. An explicit upper bound on the extinction probability function is also derived. Moreover, we establish several precise relationships between Khinchin families and random variables arising in Galton–Watson processes. We identify some of these random variables as appropriately reparameterized Khinchin families. Beyond its theoretical interest, this framework offers an efficient method for simulating probabilities associated to Galton-Watson processes by means of power series and the numerical computation of its coefficients. We believe that this perspective will be valuable to both combinatorialists and probabilists, as it provides an explicit bridge between these viewpoints.
@@ 1.1. Notations
For random variables $X$ and $Y$, the notation $X\overset{d}{=}Y$ indicates that $X$ and $Y$ are equal in distribution, that is,
@@ Acknowledgements
The author thanks Professor José L. Fernández for many valuable conversations, and also Odi Soler i Gibert for comments on an early draft of this note.
@@ 2. Khinchin families
Notice that $f(t)>0$ for each $t\in[0,R)$. We define $X_{0}\equiv 0$. Formally, the definition of $X_{0}$ is consistent with the general expression for $t\in(0,R)$, with the convention that $0^{0}=1$, meaning that $\mathbf{P}(X_{0}=0)=1$.
@@ 2.2. The classes $\mathcal{K}_{s}$ and $\mathcal{K}^{\star}$
and $Y_{t}\overset{d}{=}X_{t}+N$, where $X_{t}$ is the Khinchin family of $f$. The class $\mathcal{K}_{s}$ comprises all such shifted power series (with associated shifted Khinchin family). See [9], for further details.
@@ 3.1. The combinatorial class of rooted plane trees
We denote by $(\mathcal{G},|\cdot|)$ the combinatorial class of rooted plane trees, where the size function $|\cdot|$ gives the number of nodes in a given tree. This family of rooted trees is codified, in the language of the combinatorial classes, by means of the specification
@@ 3.3. Galton–Watson processes
with $Y_{j}$ i.i.d. copies of $Y$ and $Z_{0}=1$. The process $(Z_{n})_{n\geq 0}$ is a Galton-Watson process. The mean of the offspring distribution $m=\mathbf{E}[Y]$ determines extinction: if $m\leq 1$, extinction occurs almost surely ($q=1$); if $m>1$, the extinction probability is less than one, i.e. $q<1$. The probability $q$ is the minimal solution of the equation
@@ 4. Lagrange’s inversion formula
The following classical result describes an explicit relationship between the coefficients of the power series $\psi$ and the solution $g$, thereby providing an explicit power series representation of $g$ as a solution to Lagrange’s equation $(\star)$.
@@ 4.1. Function $t/\psi(t)$
$g(z/\psi(z))=z\,,\quad\mbox{ for }z\in\Omega_{g}=g(\mathbb{D}(0,R_{g}))\subseteq\mathbb{D}(0,R_{\psi})\,.$
@@
that is, the image of the interval $0,\rho]$ by $g$ is the interval $g([0,\tau/\psi(\tau)])=[0,\tau]$. One may verify this claim directly (using that $g$ is increasing and continuous on the closure of $\mathbb{D}(0,\rho)$), or else deduce it as a corollary of Theorem B below.
@@
    is strictly increasing on the interval $0,\tau)$
@@ 4.2. Radius of convergence of $g$
In case that $\psi\in\mathcal{K}^{\star}$ we have $R_{g}=\rho=\tau/\psi(\tau)$ .
@@
Note that when $\psi\in\mathcal{K}^{\star}$, the Khinchin family $(Z_{t})$ associated with the solution of Lagrange’s equation is naturally parametrized by $t\in[0,\rho]$, where $\rho=\tau/\psi(\tau)$. For further details about extending Khinchin families to boundary values we refer to [9].
@@ 4.3. Otter-Meir-Moon asymptotic formula
Recall that for power series $\psi(z)=\sum_{n\geq 0}b_{n}z^{n}\in\mathcal{K}$ we denote
$Q_{\psi}\triangleq\gcd\{n\geq 1:b_{n}\neq 0\}=\lim_{N\to\infty}\gcd\{1\leq n\leq N:b_{n}\neq 0\}.$
@@ 5. Lagrange’s equation and simple varieties of trees
Fix $\psi(z)=\sum_{n\geq 0}b_{n}z^{n}\in\mathcal{K}$ with radius of convergence $R_{\psi}>0$. A simple variety of trees is a family of weighted rooted plane trees whose weighted generating function is the solution of Lagrange’s equation with data $\psi$, that is, is a power series $g$ verifying the relation
@@
Here $k_{j}(a)$ counts the number of nodes in $a$ with exactly $j$ descendants. Observe that the weight $\omega$ depends on the power series $\psi$, but we omit this from the notation for simplicity. The next result is standard, see [23] for similar arguments.
@@
Let $\psi\in\mathcal{K}$ be a power series with positive radius of convergence, then
@@ 6. The parametric Galton Watson process
Given a power series $\psi\in\mathcal{K}$ and Khinchin family $(Y_{t})$, in this section we study the Galton-Watson process with offspring distribution $Y_{t}$. This construction gives rise to a family of parametric Galton-Watson processes $T_{t}$ which depends on $t\in(0,R_{\psi})$. For each $t\in(0,R_{\psi})$ the rooted random tree $T_{t}$ takes values on the Polish space $\mathbb{T}=\mathcal{G}\cup\mathcal{G}_{\infty}$.
@@ 6.1. The Galton-Watson process with offspring distribution $Y_{t}$
with the convention that $0^{0}=1$. Recall that $k_{j}(a)$ counts the number of nodes in $a$ with exactly $j$ descendants. This measure assigns probability 0 to the singletons in $\mathcal{G}_{\infty}$.
@@
with the convention that $0^{0}=1$. Here, $\omega(a)$ denotes the weight of the rooted plane tree $a\in\mathcal{G}_{n}$.
@@
these equations express that the tree has $n$ nodes and n1n-1 edges (or, equivalently, n1n-1 nodes that descend from a parent node, we don’t count the root), we obtain that
@@ 6.2. Probability generating functions
The power series $g_{t}(z)$ verifies Lagrange’s equation with data $\psi_{t}$ by definition. For any $t\in(0,R_{\psi})$, denote $f_{t}(z)={g(tz/\psi(t))}/{t}$. Notice that
@@ 6.3. Extinction probability
The probability of extinction $q(t)$ is continuous on the interval $0,R_{\psi})$.
@@
If $\psi\in\mathcal{K}\setminus\mathcal{K}^{\star}$, we have $q(t)\equiv 1$, for any $t\in[0,R_{\psi})$, which is continuous. Assume that $\psi\in\mathcal{K}^{\star}$, the function $t/\psi(t)$ is continuous for any $t\in[0,R_{\psi})$. Now, combining Theorem 4.2 with Lemma 4.1, we find that
@@
$q(t)\sim\frac{\psi(0)}{\psi(t)}=\mathbf{P}(Y_{t}=0)\,,\quad\text{ as }t\uparrow R_{\psi}\,.$
@@
Theorem 6.4 gives that
@@
  • If $R_{\psi}<\infty$, then $\lim_{t\uparrow R_{\psi}}\psi(t)<\infty$ and the probability of extinction $q(t)$ can be extended to $t=R_{\psi}$.
@@
Now we give an upper bound for the extinction probability.
@@
For any $t\in(0,R_{\psi})\setminus\{\tau\}$, combining Theorem 6.4, equation (6.6) and the identity $\psi(tq(t))=q(t)\psi(t)$ we obtain that
@@ 6.4. The uniform probability
with the convention that $0^{0}=1$. By means of $\omega$, see Proposition A above, we can build all the families of simple varieties of trees.
@@
again with the convention that $0^{0}=1$. Assume, for simplicity, that $A_{n}\neq 0$ for any $n\geq 1$, otherwise, we can restrict $n\geq 1$ to values satisfying $n\equiv 1\mod Q_{\psi}$. First, we consider the measure on $\mathcal{G}$ with weights $\omega_{\mathbf{P}}$ and $\omega$. For a rooted plane tree $a\in\mathcal{G}$ we compute its measure as follows